OpenStudy (anonymous):

solve for x: 5/x^2-1 - 2/x = 2/x+1

OpenStudy (anonymous):

Use this sustituion and carry out usual method for quadratics 1/x -> t

OpenStudy (anonymous):

so would the answer bw x=1/4 and x=-2?

OpenStudy (anonymous):

actually write the equation a bit more clearly

OpenStudy (anonymous):

can you draw it out for me? im so confused

OpenStudy (anonymous):

\[5\div((x+1)(x-1))-2/x = 2/(x+1)\] bring 2/(x+1) to lhs take LCM then crossmultiply and you get a solveable solution It is difficult to type it out

OpenStudy (anonymous):

which one of these could it be?